ac-1 - a-1c
(A times C to the negative one MINUS A to the negative one times C)

Answer by fractalier(2101)

ac^(-1) - a^(-1)c =
a/c - c/a =
a^2/ac - c^2/ac =
(a^2 - c^2) / ac